Find the value of the electric field at a distance r=10 cm from the center of a non-conducting sphere of radius R=1 cm which has an extra positive charge equal to 7 C uniformly distributed within the volume of the sphere.

1
Expert's answer
2022-03-11T08:29:26-0500

According to the shell theorem, the electric field created by the sphere is the same as the electric field of the same amount of point charge located at the center of the sphere:


"E = k\dfrac{q}{r^2} = 9\times 10^9N\cdot m/C^2\cdot \dfrac{7C}{0.1m^2} = 6.3\times 10^{12}V/m"

Answer. "6.3\times 10^{12}V/m"